OpenStudy (anonymous):

Add: 4/x^2-10x+24 + x+1/x-4

OpenStudy (anonymous):

\[\frac{ 4 }{ x ^{2}-10x+24 } + \frac{ x+1 }{ x-4 }\]

OpenStudy (anonymous):

lets factor out the \[x^2-10x+24\]

OpenStudy (anonymous):

So x(x-12)(x+2) ?

OpenStudy (anonymous):

Oh wait that would make -24 @_@.

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Or is that right?

OpenStudy (anonymous):

nope. thats wrong two numbers that multiply to +24 and add up to -10

OpenStudy (anonymous):

-6 & -4

OpenStudy (anonymous):

bingo so, the factors are: \[(x-6)(x-4)\]

OpenStudy (anonymous):

x(x-6)(x-4) right?

OpenStudy (anonymous):

now, lets write the give expression again \[ {4\over(x-6)(x-4)}+{x+1\over x-4} \]

OpenStudy (anonymous):

there are only two terms no neeed for the extra "x"

OpenStudy (anonymous):

now, just like the regular fractions, we make the denominator common

OpenStudy (anonymous):

\[ {4\over(x-6)(x-4)}+{x+1\over x-4}\times{x-6\over x-6} \]

OpenStudy (anonymous):

\[ =\frac{4+(x+1)(x-6)}{(x-6)(x-4)}\\ =\frac{4+x^2-6x+x-6}{(x-6)(x-4)}\qquad\text{I expanded the product}\\ =\frac{x^2-5x-2}{(x-6)(x-4)} \] The numerator cannot be simplified easily. so, we can leave our answer here

OpenStudy (anonymous):

Thanks so much! :)

OpenStudy (anonymous):

yw